Johnny is designing a rectangular poster to contain 24in^2 of printing with a 3-in margin at the top and bottom and a 2-in margin at each side. what overall dimensions will minimize the amount of paper used?
1
Expert's answer
2020-04-09T14:42:33-0400

Let the paper size be "y" inches in length and "x" inches in width.

The length of the printed space would be "(y-2\cdot3)" inches and width would be "(x-2\cdot2)" inches.

Johnny is designing a rectangular poster to contain "24\ in^2" of printing 


"(x-4)\cdot(y-6)=24"

Solve for "y"


"y={24\over x-4}+6"

Since the area of the paper of size "x" inches by "y" inches is "xy," let it be denoted as A


"A=x\cdot y"

Then


"A=A(x)=x({24\over x-4}+6), x>4"

Find the first derivative with respect to "x"


"A'(x)=({24x\over x-4}+6x))'={24\over x-4} -{24x\over (x-4)^2}+6"

Find the critical number(s)


"A'(x)=0=>{24\over x-4} -{24x\over (x-4)^2}+6=0"

"4(x-4)-4x+(x-4)^2=0"

"4x-16-4x+x^2-8x+16=0"


"x(x-8)=0"

Critical numbers: "x=0, x=8"

First derivative test

"If \ x<0, A'(x)>0, A(x)\ increases."

"If \ 0<x<8, A'(x)<0, A(x)\ decreases."

"If \ x>8, A'(x)>0, A(x)\ increases."

The function "A(x)" has the local maximum at "x=0."

The function "A(x)" has the local minimum at "x=8."

We consider "x>4." Hence the function "A(x)" has the absolute minimum for "x>4" at "x=8."

Find the length


"y={24\over 8-4}+6=12"

We need to use the paper "8\ in\times 12 \ in."